OpenStudy (anonymous):

Find the slope of the line passing through the given points. Then tell whether the line falls, rises, is horizontal, or is vertical. (14,-3),(4,11)

hero (hero):

Are you familiar with y = mx + b?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i know how to find the slope i just dont know how to find out whether if it falls, rises that stuff

hero (hero):

Okay, good, now, I'm going to simply plug each point into that formula: -3 = 14m + b 11 = 4m + b Are you with me so far?

OpenStudy (anonymous):

yes

hero (hero):

Okay, m is the slope, so I'm going to isolate b in both equations: b = -3 - 14m b = 11 - 4m Do you see how I got that?

OpenStudy (anonymous):

kinda

hero (hero):

It takes a little practice to understand it, but now I'm going to set b = b to get: -3 - 14m = 11 - 4m

hero (hero):

Now all I have to do is solve for m: -3 - 11 = 14m - 4m -14 = 10m -14/10 = m - 7/5 = m

hero (hero):

There we go

OpenStudy (anonymous):

ok i was wait thats not right

hero (hero):

You entered the answer as soon as I posted it, lol

OpenStudy (anonymous):

kk i got that too

OpenStudy (anonymous):

haha yeah

hero (hero):

You shouldn't enter the answer so fast :P I guess I can take my own advice in that regard.

OpenStudy (anonymous):

okay

hero (hero):

Sometimes, I make typos, so you should wait at least one minute after someone posts their answer.

OpenStudy (anonymous):

um ok

hero (hero):

My fingers are not as bright as my brain

hero (hero):

Sometimes, my fingers hit keys they shouldn't.

OpenStudy (anonymous):

oh okay its a

OpenStudy (anonymous):

all G

hero (hero):

By the way, If the slope is NEGATIVE, it FALLS If the slope is POSITIVE, it RISES

OpenStudy (anonymous):

always?

hero (hero):

If something falls, it's always going down, right? If a plane rises from the ground, it is always going up, right?

OpenStudy (anonymous):

right

hero (hero):



OpenStudy (anonymous):

ohh i see

hero (hero):

Yeah just think about a plane either taking a nose dive or lift off

OpenStudy (anonymous):

ohh okay thank u!

hero (hero):